OpenStudy (anonymous):

differentiate y=cos^-1(tan3x^4)

OpenStudy (anonymous):

Please let me first work through it. It will take one minute.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y=\cos^{-1}(\tan~(3x^4)~) }\) at first, \(\large\color{black}{ \displaystyle \cos(y)=\tan~(3x^4)~ }\)

OpenStudy (anonymous):

SO basically in derivatives for sin cos and tan, The derivative of sin x is cos x, The derivative of cos x is −sin x (note the negative sign!) and The derivative of tan x is sec2x.

OpenStudy (solomonzelman):

then use Implicit Differentiation

OpenStudy (anonymous):

Well solomon seems to have got this lol

OpenStudy (anonymous):

So the answer is cos(y)=tan(3x^4) or is there still more steps.